Q2) Ali’s budget line relating good X and good Y has intercept of 20 unit of good X and 30 units of good Y. if the price of good Y is 10, what is Ali’s income? What is the price of good Y? What is then slope of budget line
1
Expert's answer
2016-01-19T08:28:40-0500
If a budget line relating good X and good Y has intercept Qx = 20 unit of good X and Qy = 30 units of good Y and Py = $10, then the Ali’s income is:
I = Px*Qx + Py*Qy
When the budget line has the intercept Qy = 30 and Py = $10, then I = Px*0 + 30*10 = 300, so the Ali's income is I = $300.
The price of good Y can be found from the income equation in the point of interception with the X-axis:
300 = Px*20 + 10*0,
Px = $15.
The slope of budget line is:
300 = 15*Qx + 10*Qy
10Qy = 300 - 15Qx
Qy = -1.5Qx + 30, so the slope of budget line is -1.5.